# Prompt-Guided Latent Diffusion with Predictive Class Conditioning for 3D Prostate MRI Generation

### v1

@@ 3.5 Downstream Classification Task
The trained CCELLA LDM generated one synthetic image for every real training sample using the sample’s clinical text. Both the radiologist label from the source training sample (“clinical label”) and the CCELLA-generated label (“synthetic label”) were stored for each synthetic image during generation. An EfficientNet-b0 [32] was trained in each of five scenarios: 1) real data, 2) synthetic data with clinical labels, 3) synthetic data with synthetic labels, 4) real and synthetic data with clinical labels, 5) real and synthetic data with synthetic labels for synthetic images. For each scenario, the same real and synthetic data were used in training (i.e. synthetic data was not regenerated between scenarios). Models were trained across four Tesla V100 GPUs with a batch size of 16 for 16k steps at an initial learning rate of $10^{-4}$ with the AdamW optimizer and a second-order polynomial learning rate scheduler. Trained EfficientNet-b0 performances were assessed using classification accuracy on the same unseen test set as the LDM evaluation. Secondary assessment was performed using sensitivity, specificity, positive predictive value (PPV), and negative predictive value (NPV).
@@ 4.1 Performance Comparison
BASE 0.071 10.03 8.75 13.24
RADBASE 0.077 10.96 9.53 14.10
ELLABASE 0.080 11.64 10.40 15.20
@@
Figure 2: Four sample axial T2 prostate MRI images generated by CCELLA
@@ 4.3 Downstream Classification Task
Model AUC AP Accuracy Sensitivity Specificity PPV NPV % Pos
@@
Table 3 shows the performance of EfficientNet-b0 classifiers trained for identifying suspicious lesions in prostate MRI trained on any of real ($R$), synthetic with clinical label ($S_{R}$), and synthetic with synthetic label ($S_{S}$) datasets. Area under the receiver operating characteristic curve (AUC) and average precision (AP) report diagnostic performance agnostic of prediction threshold. Accuracy (Acc.), sensitivity (Sens.), specificity (Spec.), positive predictive value (PPV) and negative predictive value (NPV) were calculated at a threshold of 0.5. Augmenting real data with synthetic data for training improved all classifier metrics regardless of synthetic ground truth type, with the exception of specificity when using synthetic data with synthetic labels which remained the same. Training on synthetic data using synthetic labels as ground truth alone achieved highest AUC, AP, specificity and PPV among classifiers assessed. CCELLA generated a negative classification by default when text was missing from an image, resulting in a lower proportion of positive labels (Pos) for training for $S_{S}$. All 468 cases in the training dataset where CCELLA extracted negative classification from a positive source image occurred when clinical report text was missing. The test set had a positive case percentage of 56%.
@@ 5 Discussion
This study sought to introduce a new framework for training higher-quality synthetic prostate MRI using a limited and fixed volume dataset with limited supervised labeling. This study achieved this by proposing CCELLA and a joint loss function, in addition to reusing pretrained components where possible. Our proposed strategy achieved higher performance than baseline LDMs conditioned on nothing, on radiology label only, and on text only using ELLA. The first ablation study sought to understand the mechanism by which CCELLA achieved better performance. This ablation study trained: 1) an LDM receiving text conditioning by cross-attention using ELLA and timestep conditioning using the ground truth radiology label, 2) an LDM fitted with CCELLA but without using the proposed joint loss. CCELLA outperformed both of these variants suggesting that it is not simply the addition of a class label in the timestep embedding (DECOUPLED) nor the extension of the text adapter for timestep conditioning (L1ONLY) that results in CCELLA’s success. These results suggest that the joint training of the classifier with the text adapter encourages CCELLA to extract more clinically relevant information from text encoded by a non-clinically-trained LLM. While further investigation should be performed to confirm these findings, this ablation study suggests that pretrained LLMs outside of the medical space could be better adapted for use on medical data when trained jointly with clinical labels in lieu of adapting a pretrained, non-medical LLM without additional labels or training a new LLM on clinical information.
@@
Sample images from the CCELLA-trained LDM can be seen in Figure 2. Qualitatively, these images exhibit natural variance in the sizes and shapes of the prostate while consistently depicting the different anatomical regions within the prostate. Some images display differences in brightness or smoothness of the image which can give a slightly less “realistic” appearance despite the anatomical accuracy and the utility of these images in augmenting the dataset of a downstream task. Thus, there exist possible research opportunities in developing strategies for the measuring and improving the consistency of image quality during synthetic medical image generation with diffusion models.
@@
Finally, this study trained downstream classifier models for prostate cancer prediction from MRI using real data and synthetic data generated by a CCELLA-trained LDM. The $R+S_{R}$ and $R+S_{S}$ classifiers performed equally or better than the $R$ classifier on all metrics, with $R+S_{R}$ outperforming $R$ on all confusion matrix metrics. These results validate the hypothesis that synthetic data can be used to successfully augment the training dataset for a classifier model and aligns with recent literature that generate ground truths alongside data samples. $S_{S}$ also performed comparably to $R$ with higher AUC and AP but lower accuracy, suggesting that the CCELLA-generated synthetic data may be as useful as real data when training downstream machine learning tasks after tuning the classifier threshold. These results are promising, though it is important to note that the synthetic data was generated from the same distribution as the real data. Future work should seek to investigate these results further by generating synthetic data for training with real data of different distribution.
@@
Interestingly and in contrast to expectation, the $S_{S}$ classifier, trained only on the same synthetic data as $S_{R}$ but with ground truth labeling sourced from CCELLA’s classifier head instead of the clinical label accompanying the source image, outperformed not only the $S_{R}$ classifier in all metrics except sensitivity, but also the $R$ classifier in multiple metrics (AUC, AP, specificity, PPV). Seemingly paradoxically, augmenting the real dataset with $S_{R}$ resulted in better performance than augmenting with $S_{S}$ in all metrics except AP and sensitivity where they were equal. This likely occurred for three reasons. Firstly, only 3701 of the 5071 training samples had clinical report text. Because null text reports generated negative class predictions by default, the label distribution for synthetic labeling was different from from the real label distribution which would have affected the model training. It therefore makes retrospective sense that $R+S_{S}$, having trained on a dataset with lower positive case percentage than $R+S_{R}$ (43% versus 48%), would exhibit more difficulty in adapting to a test set with higher positive case percentage (56%), resulting in lower accuracy (0.71 versus 0.74). Secondly and related to the first, the training cases that were missing radiology text belonged to both positive and negative classes. Thus even though the CCELLA classifier head was not trained on these cases, the LDM would still have learned to generate both positive and negative images when presented with a null prompt. This in-turn could have affected the performance of classifiers trained with synthetic data. Thirdly, the CCELLA LDM did not receive the real image as input during inference. Consequently, even though the ground-truth radiology classification was based on the radiologist re-interpretation of the full image under the most recent prostate MRI reporting standard instead of simply extracting the classification in the clinical text, the use of the ground truth re-classification may have lost its meaning in favor of what the LLM interprets from the clinical text. This could have resulted in the synthetic images aligning well with the text interpretation and CCELLA class extraction but not necessarily with the radiologist re-classification of the real image corresponding to that text, which could explain why $S_{R}$ exhibited poorer performance. Thus, while the same train and test datasets used for LDM training were used for the downstream classifier task for the sake of consistency, future work should seek to further explore the difference between using real versus synthetic labels for synthetic images for the purpose of dataset augmentation when training a model for a medical task on data.
@@
This study has a few limitations, each of which present an opportunity for future work. Firstly, this study emphasized data from a single institution, for a single modality in a single pathology. While the pathology and institution restrictions were intentional to the study design in developing a limited-data LDM training framework, future work should seek to adapt this framework to multi-institutional data, as well as images of different modality or pathology to assess method generalizability. Additionally, prostate MRI is clinically assessed in a minimum of three, and preferably four different MRI sequences [33, 34]. Consequently, portions of the clinical text used for conditioning would have referred to MRI sequences that the LDM was neither trained on nor asked to generate. While the use of one sequence alone in this study was chosen to focus primarily on the adapter innovation and data-limited framework design instead of the additional challenge of multi-sequence generation, future work should seek to extend data-limited LDM training strategies for optimal multi-sequence generation. As well, this study included medical images from public datasets that did not have accompanying text reports for the purpose of gathering more training samples to train the denoising U-Net and the downstream classifier models. While outside the scope of this study, this prompts two opportunities for future work. One opportunity is to investigate how the ratio of conditioned to unconditioned data samples affect model training and whether there is a critical ratio below which training with conditioning becomes unstable or negatively affects training. Another opportunity is to explore synthetic clinical text report generation for these data samples, whereby either a rule-based algorithm or LLM informed with the radiologist classification and image metadata provided in the public dataset would generate text reports for use in conditioning the LDM during training and inference. Finally, while the intent of this study aligns with a growing body of literature seeking to use LDMs to generate synthetic medical data for the purpose of mitigating data scarcity challenges within machine learning for medical imaging applications, recent literature has also shown that LDMs can occasionally memorize and reproduce their training samples on inference [35]. While outside the scope of this study, future work investigating and improving upon differential privacy within medical imaging LDMs would be valuable.
@@ 6 Conclusion
In summary, this study presents CCELLA, a novel adapter for simultaneous alignment of non-medical LLMs to the medical image space and pathology extraction from medical text for LDM conditioning, as well as a CCELLA-centric framework for medical image LDM training given limited data volume and annotation. Our proposed method outperforms comparative baseline models when generating synthetic prostate MRI data and successfully augments the training dataset of a downstream classifier model for improved cancer prediction performance over training on real data alone. The results of this study suggest that CCELLA can successfully adapt pretrained, non-medical LLMs for use in conditioning LDMs for medical image generation, and can successfully train a 3D medical image LDM using data generated during clinical routine. Future work should focus on expanding this framework to multi-sequence image generation, validating the framework in a different medical imaging and pathology space, and exploring differential privacy within medical imaging LDMs.

### v2

@@ 3.4 Downstream Classification Task
The trained CCELLA LDM generated one synthetic image for every real training sample using the sample’s clinical text. Both the radiologist label from the source training sample (“clinical label”) and the CCELLA-generated label (“synthetic label”) were stored for each synthetic image during generation. An EfficientNet-b0 [32] was trained in each of five scenarios: 1) real data, 2) synthetic data with clinical labels, 3) synthetic data with synthetic labels, 4) real and synthetic data with clinical labels, 5) real and synthetic data with synthetic labels for synthetic images. For each scenario, the same real and synthetic data were used in training (i.e. synthetic data was not regenerated between scenarios).
Training details: All synthetic images were resampled to the original voxel size of their corresponding real image using Lanczos interpolation [28]. All images were linearly rescaled from 0%-98% intensity with clipping then padded to a $128mm\times 128mm\times 96mm$ field of view. Images were augmented during training with random flips, scaling, translations, and rotations. All classifiers were trained on four Tesla V100 GPUs with a batch size of 16 for 16k steps at an initial learning rate of $10^{-4}$ with the AdamW optimizer and a second-order polynomial learning rate scheduler.
Evaluation: Trained EfficientNet-b0 performances were assessed using the area under the receiver operating characteristic curve (AUC), average precision (AP), and classification accuracy on the same unseen test set as the LDM evaluation. Secondary assessment was performed using sensitivity, specificity, positive predictive value (PPV), and negative predictive value (NPV).
@@ 4.1 Performance Comparison
MAISI 0.071 10.03 8.75 13.24
RADMAISI 0.077 10.96 9.53 14.10
ELLAMAISI 0.080 11.64 10.40 15.20
@@
Figure 2: Four real axial T2 prostate MRI images and the synthetic MRI images generated by ELLAMAISI and CCELLA conditioned on the same corresponding report text.
@@ 4.3 Downstream Classification Task
Model AUC AP Acc. Sens. Spec. PPV NPV % Pos
@@
Table 3 shows the performance of EfficientNet-b0 classifiers trained for identifying suspicious lesions in prostate MRI trained on any of real ($R$), synthetic with clinical label ($S_{R}$), and synthetic with synthetic label ($S_{S}$) datasets. AUC and AP report diagnostic performance agnostic of prediction threshold. Accuracy (Acc.), sensitivity (Sens.), specificity (Spec.), positive predictive value (PPV) and negative predictive value (NPV) were calculated at a decision threshold of 0.5. Augmenting real data with synthetic data for training improved all classifier metrics regardless of synthetic ground truth type, with the exception of specificity when using synthetic data with synthetic labels which remained the same. Training on synthetic data using synthetic labels as ground truth alone achieved highest AUC, AP, specificity and PPV among classifiers assessed. CCELLA generated a negative classification by default when text was missing from an image, resulting in a lower proportion of positive labels (Pos) for training for $S_{S}$. All 468 cases in the training dataset where CCELLA extracted negative classification from a positive source image occurred when clinical report text was missing. The test set had a positive case percentage of 56%.
@@ 5 Discussion
This study sought to introduce a new framework for training higher-quality synthetic prostate MRI using a limited and fixed volume dataset with limited supervised labeling. This study achieved this by proposing CCELLA and a joint loss function, in addition to reusing pretrained components where possible. Our proposed strategy achieved higher performance than baseline LDMs conditioned on nothing, on radiology label only, and on text only using ELLA. The first ablation study sought to understand the mechanism by which CCELLA achieved better performance. This ablation study trained: 1) an LDM receiving text conditioning by cross-attention using ELLA and timestep conditioning using the ground truth radiology label, 2) an LDM fitted with CCELLA but without using the proposed joint loss. CCELLA outperformed both of these variants, confirming the primary study hypothesis that it is not simply the addition of a class label in the timestep embedding (DECOUPLED) nor the extension of the text adapter for timestep conditioning (L1ONLY) that results in CCELLA’s success. In combination, these ablation results indicate that the joint training of the classifier with the text adapter encourages CCELLA to extract more clinically relevant information from text encoded by a non-clinically-trained LLM. These ablation results also suggest that pretrained, non-medical LLMs could be better adapted for medical text processing when trained jointly with clinical labels. Further investigation should be performed to confirm these findings in different medical domains.
@@
Figure 2 shows four sample real images alongside the synthetic images generated by CCELLA and ELLAMAISI using the same corresponding text prompts. Qualitatively, the CCELLA images exhibit natural variance in the sizes and shapes of the prostate while consistently depicting the different anatomical regions within the prostate. CCELLA also did not struggle with image brightness and contrast to the same degree as ELLAMAISI for the same text prompts. ELLAMAISI exhibited occasionally difficulty in defining prostate regions, showing a clear artificial smoothness in Case 4 of Figure 2. The qualitative improvements of CCELLA over ELLAMAISI reaffirm the design choice for the CCELLA adapter to use class conditioning as a directly trainable output.
@@
Finally, this study trained downstream classifier models for prostate cancer prediction from MRI using real data and synthetic data generated by a CCELLA-trained LDM. The $R+S_{R}$ and $R+S_{S}$ classifiers performed equally or better than the $R$ classifier on all metrics, with $R+S_{R}$ outperforming $R$ on all confusion matrix metrics. These results validate the hypothesis that synthetic data can be used to successfully augment the training dataset for a classifier model and aligns with recent literature that generate ground truths alongside data samples. $S_{S}$ also performed comparably to $R$ with higher AUC and AP but lower accuracy, suggesting that the CCELLA-generated synthetic data may be as useful as real data when training downstream machine learning tasks after classifier threshold tuning. These results are promising, though it is important to note that the synthetic data was generated from the same distribution as the real data. Future work should seek to investigate these results further by generating synthetic data for training with real data of different distribution.
@@
Interestingly and in contrast to expectation, the $S_{S}$ classifier outperformed not only the $S_{R}$ classifier in all metrics except sensitivity, but also the $R$ classifier in multiple metrics (AUC, AP, specificity, PPV). Seemingly paradoxically, augmenting the real dataset with synthetic data in $R+S_{R}$ resulted in better performance than augmenting in $R+S_{S}$ in all metrics except AP and sensitivity where they were equal. This likely occurred for three reasons. Firstly, only 3701 of the 5071 training samples had clinical report text. Because null text reports generated negative class predictions by default, adding synthetic data with synthetic labeling in $R+S_{S}$ resulted in a lower positive case percentage than using real labels in $R+S_{R}$ (43% versus 48%). It therefore makes retrospective sense that $R+S_{S}$ would exhibit more difficulty than $R+S_{R}$ in adapting to a test set with higher positive case percentage (56%), resulting in lower accuracy (0.71 versus 0.74). Secondly and related to the first, the training cases without radiology text belonged to both positive and negative classes. Even though the joint classifier loss was removed for these cases, the LDM would still have learned to generate both positive and negative images when presented with a null prompt. This in-turn could have affected the performance of classifiers trained with synthetic data. Thirdly, the CCELLA LDM did not receive the real image as input during inference. However, the ground-truth radiology classification was based on the radiologist re-interpretation of the full image under the most recent prostate MRI reporting standard instead of simply extracting the classification in the clinical text. Consequently, the ground truth re-classification may have lost its meaning in favor of what the LLM interpreted from the clinical text. This could have resulted in the synthetic images aligning well with the text interpretation and CCELLA class extraction but not necessarily with the radiologist re-classification of the real image corresponding to that text, which could explain why $S_{R}$ exhibited poorer performance. Thus, while the same train and test datasets used for LDM training were used for the downstream classifier task for the sake of consistency, future work should seek to further explore the difference between using real versus synthetic labels for synthetic images for the purpose of dataset augmentation when training a model for a medical task on data.
@@
This study has a few limitations, each of which present an opportunity for future work. Firstly, this study emphasized data from a single institution, for a single modality in a single pathology. While the pathology and institution restrictions were intentional to the study design in developing a limited-data LDM training framework, future work should seek to assess method generalizability by adapting this framework: a) to multi-institutional data, b) to images of different modality or pathology. Additionally, prostate MRI is clinically assessed in a minimum of three, and preferably four different MRI sequences [33, 34]. Consequently, portions of the clinical text used for conditioning would have referred to MRI sequences that the LDM was neither trained on nor asked to generate. The use of one sequence alone in this study was chosen to focus primarily on the adapter innovation and data-limited framework design instead of the additional challenge of multi-sequence generation. Nonetheless, future work should seek to extend data-limited LDM training strategies for optimal multi-sequence generation. As well, this study included medical images from public datasets that did not have accompanying text reports for the purpose of gathering more training samples to train the denoising U-Net and the downstream classifier models. While outside the scope of this study, this prompts two opportunities for future work. One opportunity is to investigate how the ratio of conditioned to unconditioned data samples affect model training and whether there is a critical ratio below which training with conditioning becomes unstable or negatively affects training. Another opportunity is to explore synthetic clinical text report generation for these data samples. Finally, the intent of this study aligns with a growing body of literature seeking to use LDMs to generate synthetic medical data for the purpose of mitigating data scarcity challenges within machine learning for medical imaging applications. However, recent literature has also shown that LDMs can occasionally memorize and reproduce their training samples on inference [35]. While outside the scope of this study, future work investigating and improving upon differential privacy within medical imaging LDMs would be valuable.
@@ 6 Conclusion
In summary, this study presents CCELLA: a novel adapter for simultaneous alignment of non-medical LLMs to the medical image space and radiology label extraction from medical text for LDM conditioning. This study also presents a CCELLA-centric framework for medical image LDM training given limited data volume and annotation. Using limited data and minimal human annotation, our proposed framework outperforms a recent medical LDM foundation model and successfully augments the training dataset of a downstream classifier model for improved cancer prediction performance over training on real data alone. The results of this study suggest that CCELLA can successfully adapt pretrained, non-medical LLMs for use in conditioning LDMs for medical image generation, and can successfully train a 3D medical image LDM using data generated during clinical routine. Future work should focus on expanding this framework to multi-sequence image generation, validating the framework in a different medical imaging and pathology space, and exploring differential privacy within medical imaging LDMs.